OpenStudy (anonymous):

@xMissAlyCatx @alekos @ganeshie8 can someone help?

OpenStudy (anonymous):

A recipe calls for 2 cups of flower to make 24 cookies. write and solve a proportion to find how much flour is needed to make 54 cookies. 2/24=x/54; x=4 1/2 cups 2/54=x/24; x=8/9 cups 2/54=24/x; x=27 cups 2/24=54/x; x=648 cups

OpenStudy (anonymous):

@wes_is_here

OpenStudy (thread252):

I think it is A how bout you?

OpenStudy (anonymous):

i dont know its so confusing

OpenStudy (thread252):

IKR

OpenStudy (thread252):

that is my best guess, there is nothing else i can do

OpenStudy (wes_is_here):

ummm

OpenStudy (anonymous):

can you explain why you think its a at least

OpenStudy (thread252):

IDK my instinct is telling that one is right

Nnesha (nnesha):

\[\rm \frac{ how ~cups~of~flower }{ number~of~total~ cookie}=\frac{ how ~cups~of~flower }{ number~of~total~ cookie}\]

OpenStudy (anonymous):

i dont want to be rude but i'd rather go for an educated guess instead of instict

Nnesha (nnesha):

` for 2 cups of flower to make 24 cookies `and we have to find how many flour (cups of flour = x) we need to make 54 cookie so plugin

OpenStudy (anonymous):

so a or b?

Nnesha (nnesha):

2 cups of flower to make how many cookies ? 24 or 54 ?

OpenStudy (anonymous):

its A

OpenStudy (thread252):

so I was right!!!!!!! YAYAYAYAYYAYA

OpenStudy (anonymous):

dont push it lol

Nnesha (nnesha):

\[\rm \frac{ 2 cups }{ 24 cookie }=\frac{ x }{ 54cookie }\] x= number of cups f flower we need to make 54 cookie

Nnesha (nnesha):

now cross multiply to solve for x

OpenStudy (anonymous):

could you help with another?

OpenStudy (anonymous):

x= 4 1/2 cups

Nnesha (nnesha):

i'm sorry but i have to go ...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyone else?

OpenStudy (thread252):

no one is here to help

OpenStudy (thread252):

I guess